OpenStudy (ttp):

Simplify twenty-five to the negative ninth divided by the quantity twenty-fifth squared. one divided by the quantity twenty-five to the seventh one divided by the quantity twenty-five to the eleventh 25 to the 7th power 25 to the 11th power

OpenStudy (anonymous):

this is the problem right: \[\frac{25^{-9}}{25^2}\]

OpenStudy (anonymous):

we divide the same base with some exponent, so we can also dubtract the exponents from each other -9-2 = -11

OpenStudy (ttp):

So it's either b or c? Which one:) ?

OpenStudy (anonymous):

"twenty-fifth squared" that's is not 25^2. what is it ?

OpenStudy (anonymous):

\[(\frac{20}{5})^2 ??\]

OpenStudy (ttp):

I'm confused :/

OpenStudy (anonymous):

"twenty-fifth squared"

OpenStudy (ttp):

?

OpenStudy (ttp):

So it's not either B or D?

OpenStudy (anonymous):

I m unsure what that would be written in numbers

OpenStudy (anonymous):

is twenty-fifth 20/5 ??

OpenStudy (ttp):

No no. Twenty fifth is 25

OpenStudy (anonymous):

oh ok

OpenStudy (ttp):

yea. Answer?

OpenStudy (anonymous):

one divided by the quantity twenty-five to the eleventh

OpenStudy (anonymous):

which is the same as 25^-11

OpenStudy (ttp):

So b?

OpenStudy (anonymous):

25^-11 = 1/25^11 right ?

OpenStudy (ttp):

Yesum, Thanks hun(: